A car of mass m traveling at speed v crashes into the rear of a truck of mass 2m that is at rest and in neutral at an intersection. If the collision is perfectly inelastic, what is the speed of the combined car and truck after the collision?
1
Expert's answer
2020-11-25T10:53:45-0500

according to momentum conservation law, sum of momentums of two cars is equal to momentum of system of two cars moving together after collision.

"m_{1}\times V_{1} + m_{2} \times V_{2}=m_{3}\times V_{3}"

"m\times V+2\times m \times0=3\times m \times V_{combined}"

"V_{combined}=\frac{m\times V}{3\times m}=\frac{V}{3}"